Write standard form of the equation of line through pair of points (3,1) and (-7,7). Write euqation using only integer coefficients.
Answer by Alan3354(69443)

Write standard form of the equation of line through pair of points (3,1) and (-7,7). Write euqation using only integer coefficients.
--------------
|+x +y +1|
|+3 +1 +1| = 0
|-7 +7 +1|
-----------
x*(1-7) - y*(3+7) + 1*(21+7) = 0
-6x - 10y = -28
3x + 5y = 14